5(t+1)+8= 3(t-2)+6
Found 3 solutions by Blue_Skies, rfadrogane, MathLover1:
Answer by Blue_Skies(5)

OK! This question has to first be solved by using Distributive Property. Like so:
5(t+1)+8= 3(t-2)+6
5t + 5 + 8 = 3t - 6 + 6
Then simplify! :)
5t + 13 = 3t
13 = 3t - 5t
13 = -2t
t = 13/-2
t = -6.5
I hope I helped!

Answer by rfadrogane(214)

5(t+1)+8= 3(t-2)+6
Soln;
5(t+1)+8= 3(t-2)+6 , distribute
5t+5+8=3t-6+6 , combine like terms
5t+13=3t
(5-3)t=-13
2t=-13, divide both side by 2
t = -13/2 ---answer

Answer by MathLover1(6634)